OpenStudy (anonymous):

Perform the indicated operation and if possible simplify. -7/120 - 1/3

OpenStudy (anonymous):

its (-7/120)-(1/3) is it?

OpenStudy (anonymous):

not according to the problem on my quiz.

OpenStudy (callisto):

Is the question \(-\frac{7}{120} - \frac{1}{3}\)?

OpenStudy (anonymous):

I'm sorry, I wrote it wrong...it's supposed to be -712 - 1/3

OpenStudy (anonymous):

-7/12 - 1/3

OpenStudy (callisto):

So the question is\[-\frac{7}{12} - \frac{1}{3}\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-1/3= -4/12

OpenStudy (callisto):

\[-\frac{7}{12}-\frac{1}{3}\]\[=-\frac{7}{12}-\frac{1(4)}{3(4)}\]\[=-\frac{7}{12}-\frac{4}{12}\]\[=\frac{-7-4}{12}\]\[ = ...?\]

OpenStudy (anonymous):

-4/12?

OpenStudy (anonymous):

What is: -7 - 4 = ??

OpenStudy (anonymous):

-11

OpenStudy (callisto):

So, you'll get -11 as your numerator... Can you get the answer now??

OpenStudy (anonymous):

Write the entire..

OpenStudy (anonymous):

Yes ok, thanks guys

OpenStudy (anonymous):

Welcome.. Try to make the denominator same in terms of adding or subtraction of fractions..

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

Sorry: Try to make the denominator same in case of addition and subtraction of fractions..